Grade 12AIEEE Entrance Exam

a particle of radius r=21/22m with constant speed 1m/s find magnitude of average velocity in 2 sec

rushikeshbhavthankar
11 Years agoGrade 12

1 Answer

Sumit Majumdar
11 Years ago
Dear student,
The time period of the particle would be given by:
T=\frac{2\pi r}{v}=6 sec
So, in 3 sec, the displacement of the particle would be equal to
\frac{21}{11} m.
Hence, in 2 sec, the displacement of the particle would be equal to \frac{21}{11} \times \frac{2}{3}=\frac{14}{11} m
hence, the required velocity would be equal to:
\frac{7}{11} m/sec.
Regards
Sumit

question mark
How to join Illuminatti occult for money 💰 ritual +2347045790756!!
 
+2347045790756 #SECRET #OCCULT
It's encourageable to see an individual trying to join us in pursuit of happiness and luxury, But you must also put things into consideration thereby knowing it's takes a fortune to achieve entrance into the Grand sacred Temple of ALGON KORAH BROTHERHOOD OCCULT.
-------------- +2347045790756 ----------------
We believe everyone has the potential to become whatever they chooses but often times 70% of Humanity have failed in the race of life thereby not been able to achieve anything here on earth, maybe because of FEAR & ANXIETY.
We don't force anyone to become a member here but of what benefit are you living a poor life, without hope for better opportunities, without hope of making your family proud, without making your children proud, without making yourself proud as an individual.
The 7 Occult SECRETS  will make you to understand that there are no rituals without danger and effects but at the same time anything that has danger and effect will definitely have Guidelines and caution, So when the guidelines and caution is given then you are now left with your school of thought which permits you to commemorate on the caution and Guidelines.
-------------- +2347045790756 -----------------
Though the days of human rituals has been condemned and abolished but it's also known that ritual without sophisticated materials and mantles can never Pierce the heart of the spiritual world and it then becomes useless thereby causing harm and social disorder which might possibly ones Life.
Here in ALGON KORAH Brotherhood occult we don't settle for less, every ritual items & mantles required for ritual is been purchased no matter how much it takes.
So when you think about becoming a member of this Occult have it in mind that it takes a great sacrifice to Pierce the heart of the spiritual world in order for an individual to acquire spiritual wealth and luxury, Call now for enquiries +2347045790756
Many atimes has been defrauded in their pursuit for spiritual wealth thereby going back to the dust where they started from, why will an individual attempt joining Illuminati, Freemasons confraternity, odd-fellow. All this above mentioned confraternities they belong to different monopolies of Life which never grants you the ability to join unless you're rich and famous for them to get more glory.
So anyone telling you to come join this above mentioned confraternities will only defraud you and make you toodle in darkness, and make you an object of mediocrity in the society.
Before you can be accepted as a member here you must pass through spiritual confirmation to know if your DESTINY will match the secret heart of this occult because only those their destiny matches the secret heart of this occult can be accepted to become a member here.
Membership into the sacred heart of this brotherhood is free but you're taking responsibility of every ritual items and mantles required for your initiation rituals, Call now for enquiries +2347045790756
A certified member of this brotherhood is not allowed to introduce a new member in order for the "7 OCCULT SECRETS" of this brotherhood to be safe and not mentioned to a mediocre.
That's why we're here online for your searching eg.
 
I_WANT_TO_JOIN_OCCULT_FOR_MINEY_RITUAL
 
HOW_TO_JOIN_OCCULT_FOR_MONEY_RITUAL
 
I WANT TO JOIN OCCULT ILLUMINATI
I WANT TO JOIN REAL OCCULT TO BE RICH
WHERE TO JOIN OCCULT FOR MONEY RITUAL
I WANT TO BE RICH
I WANT TO BE FAMOUS
 
WE ARE NOW HERE FOR YOU, TO GUIDE YOU THROUGH THE ULTIMATE PART OF WEALTH POWER PROTECTION AND EVERYTHING LUXURY CAN BRING HERE ON EARTH JOIN US TODAY TO GIVE YOUR LIFE A MEANING.
 
CONTACT SPIRITUAL GRANDMASTER FOR ENQUIRIES +2347045790756
 
-------------------TEMPLE-MEDIA---------------
 
 
aieee entrance exam

Last Activity: 4 Years ago

question mark
any network theorem doesn't obeys ohms law or all theorems obeys ohms law
 
aieee entrance exam
0 Answer Available

Last Activity: 4 Years ago